In Europe, approximately 4% of people are red-haired, mostly due to a recessive mutation in the melanocortin 1 receptor (MC1R). If the population is in Hardy-Weinberg equilibrium, what is the likelihood that a non-redheaded European is a carrier for the MC1Rr allele? (For the purpose of this problem, assume all redheads and carriers are mutant for the same MC1R allele).

A

Option A: 22%

B

Option B: 33%  

C

Option C: 11%

D

Option D: 44%  
Answer #1

Let, Q= recessive

P= dominant

(Q)2 = 4/100

  Q = 0.2

Since, the population is in Hardy-Weinberg equilibrium

P + Q = 1

P = 1 - 0.2

P = 0.8

Now, total population of non-redheaded European is given by (P2 + 2PQ)

So, P2 + 2PQ = (0.8)2 + 2*0.8*0.2

= 0.64 + 0.32

= 0.96

Now, the people in heterozygous condition is the carrier (2PQ)

So, the likelihood that a non-redheaded European is a carrier for the MC1Rr allele is (0.32/ 0.96)*100 = 33.33%

Option B is correct